Alice has 20 cm of wire. She makes a rectangular shape with the
wire. The shape has the largest possible area. What is the length and
width of the shape ALice made? Explain how you worked it out.
1
Expert's answer
2019-04-16T10:10:59-0400

"2x+2y=20\:\;->\;\;y=10-x\\
S=xy=x(10-x)\\
\frac{dS}{dx}=10-2x\\
\frac{dS}{dx}=0\;\;->\;\;x=5,\;\;y=10-x=5.\\
Point\;(5,5)\;is\ the\ maximum\ point\ because\;\;\frac{d^2S}{dx^2}=-2<0"